Bal Jan 1      1,000 @ $6 = 500 units                              Purchases:

Purchase:                                                                              Jan 26              2,500 @ $16 = 500 units

Jan 6            2,000 @ $12                                                    Sales:

Sales:                                                                                      Jan 31              (2,000)

Jan 7             (2,500)

Ending Inventory 500 units @ $6 = $3,000                      Ending Inventory 500 units @ $16 = $8,000

Assuming that Rich maintains perpetual inventory records, what should be the inventory at Jan 31, using the moving-average inventory method?
Answer #1

Under Perpetual inventory moving average inventory method ,Average cost per unit is calculated at the time of every purchase and then inventory is sold at that average cost.

Average cost per unit =cost of goods available for sales /units available for sale

MOVING AVERAGE PERPETUAL INVENTORY METHOD
PURCHASE COST OF GOODS SOLD INVENTORY AT END
Quantity unit cost Total Quantity unit cost Total Quantity unit cost Total
Beginning 1000 6 6000
Jan 6 2000 12 24000 1000+2000=3000

[6000+24000]/3000

$ 10 per unit

30000
Jan 7 2500 10 25000 3000-2500=500 10 5000
Jan 26 2500 16 40000 2500+500= 3000

[5000+40000]/3000

=45000/3000

=$ 15 per unit

45000
Jan 31 2000 15 30000 3000-2000=1000 15 15000

Inventory Balance at Jan 31 = $ 15000

                    [ 1000 Units at average cost of $ 15 per unit = $ 15000]